# Multi-scale Graph Autoregressive Modeling: Molecular Property Prediction via Next Token Prediction

### v2

@@ 3.2.1 Autoregressive Pre-training via NTP
CamS-LLaMA adopts a standard LLaMA-style decoder (Touvron et al., 2023); in pre-training, every CamS token list is treated as an ordinary token sequence in the shared vocabulary $\Sigma$ and is consumed by the backbone under the usual causal mask, with no graph-specific architectural modification. Concretely, each molecule yields five CamS sequence views $\{X_{1},X_{7},X_{27},X_{67},X\}$: four single-scale BFS sequences tokenized with $\Sigma_{1\mathrm{K}}/\Sigma_{7\mathrm{K}}/\Sigma_{27\mathrm{K}}/\Sigma_{67\mathrm{K}}$, plus the concatenated multi-scale sequence $X$ in Eq. (1); we treat each view as one NTP training instance (details in Appendix A.3). Given any view $X=(x_{1},\ldots,x_{T})$, token embeddings form $\mathbf{H}^{(0)}$, which is transformed by $L$ causal self-attention blocks into $\mathbf{H}^{(L)}$; a tied output projection then yields $p_{\theta}(x_{t+1}\mid x_{\leq t})$.
Importantly, CamS changes what the model can condition on under the same NTP objective by reshaping the prefix context through fine-to-coarse concatenation (Eq. (1)). Let $\mathcal{I}_{\mathrm{tok}}\subseteq\{1,\ldots,T-1\}$ be valid prediction positions in $X$ (excluding [BOS], [CONCAT], [EOS]), and $x^{\star}_{t+1}$ be the ground-truth next token. The NTP loss is
$\mathcal{L}_{\mathrm{NTP}}(\theta;X)=-\sum_{t\in\mathcal{I}_{\mathrm{tok}}}\log p_{\theta}(x^{\star}_{t+1}\mid x_{\leq t}).$ (2)
Because $X$ concatenates BFS sequences from fine to coarse scales (Eq. (1)), the conditioning prefix $x_{\leq t}$ has an explicit hierarchy: for $t$ in $X_{67}$ (highest-merge segment), the model conditions on all finer-scale tokens $(X_{1},X_{7},X_{27})$ plus the coarse prefix. Thus, coarse-scale prediction errors propagate through attention to both coarse predecessors and the entire fine-scale prefix, providing a fine-to-coarse causal credit assignment under a fully standard decoder.
On the pre-training corpus $\mathcal{D}_{\mathrm{pre}}$, we minimize
$\min_{\theta}\;\mathbb{E}_{X\sim\mathcal{D}_{\mathrm{pre}}}\big[\mathcal{L}_{\mathrm{NTP}}(\theta;X)\big],$ (3)
learning general-purpose molecular representations purely from CamS sequences, without any downstream labels or extra descriptors.
3.2.2 Fine-tuning with Fingerprint Injection
Many molecular pre-training methods inject rich handcrafted descriptors early (e.g., during pre-training) (Li et al., 2023). In contrast, we keep CamS as the primary representation and inject only one widely-used structural prior–a fingerprint–only during fine-tuning, to make the contribution of CamS representations more attributable and controlled.
We treat the fingerprint as a virtual global token. For each molecule, a fingerprint vector
$\mathbf{f}\in\mathbb{R}^{D_{\mathrm{fp}}},$ (4)
is projected to the model hidden size:
$\mathbf{f}^{\prime}=W_{\mathrm{fp}}\mathbf{f},\quad W_{\mathrm{fp}}\in\mathbb{R}^{H\times D_{\mathrm{fp}}},$ (5)
and prepended to token embeddings:
$\mathbf{H}^{(0)}_{\mathrm{ft}}=[\,\mathbf{f}^{\prime};\mathbf{e}_{1};\dots;\mathbf{e}_{T}\,]\in\mathbb{R}^{(T+1)\times H}.$ (6)
With a modified causal mask, all positions can attend to this global token. After $L$ layers:
$\mathbf{H}^{(L)}_{\mathrm{ft}}=[\,\mathbf{h}^{(L)}_{0};\mathbf{h}^{(L)}_{1};\dots;\mathbf{h}^{(L)}_{T}\,$]. (7)
We use the global token representation $\mathbf{u}^{(i)}=\mathbf{h}^{(L)}_{0}$ for prediction (a decoder-only analogue of [CLS]).
For classification:
$\begin{gathered}\hat{\mathbf{y}}^{(i)}=\mathrm{softmax}\!\big(W_{2}\,\sigma(W_{1}\mathbf{u}^{(i)})\big),\\ \mathcal{L}_{\mathrm{cls}}(\theta,\phi)=-\frac{1}{B}\sum_{i=1}^{B}\log\hat{\mathbf{y}}^{(i)}_{\,y^{(i)}}.\end{gathered}$ (8)
For regression:
$\begin{gathered}\hat{y}^{(i)}=W_{2}\,\sigma(W_{1}\mathbf{u}^{(i)}),\\ \mathcal{L}_{\mathrm{reg}}(\theta,\phi)=\frac{1}{B}\sum_{i=1}^{B}\big(\hat{y}^{(i)}-y^{(i)}\big)^{2}.\end{gathered}$ (9)
Fine-tuning minimizes
$\min_{\theta,\phi}\;\mathbb{E}_{(X,\mathbf{f},y)\sim\mathcal{D}_{\mathrm{task}}}\big[\mathcal{L}_{\mathrm{cls}/\mathrm{reg}}(\theta,\phi;X,\mathbf{f},y)\big].$ (10)
This keeps pre-training purely sequence/structure-driven, while using fingerprints as a lightweight and controlled prior only when optimizing for downstream targets.
3.3 CamS Representation vs. Graph Transformer
@@ 3.2.2 Fine-tuning with Fingerprint Injection
We analyze CamS-LLaMA and graph transformer-style models (e.g., KPGT) within a unified framework of Token-Level Graph-Structured Deep Feature Construction. We show that CamS yields (1) a denser and multi-view direct supervision signal and (2) a hierarchical inductive bias that are mathematically distinct from standard masked-node prediction (MNP). Detailed derivations and supporting discussions are provided in Appendix B.
3.3.1 Unified Framework.
@@
Let $G=(V,E)$ be a molecular graph. Both paradigms construct a token graph $\mathcal{G}_{\mathrm{tok}}=(\mathcal{V}_{\mathrm{tok}},\mathcal{E}_{\mathrm{tok}})$ and perform $L$ layers of propagation:
@@
where $\mathbf{H}^{(l)}$ are token embeddings, $\mathbf{q}_{i},\mathbf{k}_{j}$ are projections of $\mathbf{H}^{(l)}$, and $\psi_{ij}$ encodes the structural bias or causal mask.
Graphformer (Flat/Static Bias). $\mathcal{V}_{\text{tok}}$ corresponds to single-scale atoms or triplets. $\psi_{ij}$ encodes static structural biases (e.g., shortest-path-distance bias injected into attention) independent of token content, yielding an isotropic receptive field at a given graph distance (Ying et al., 2021).
CamS-LLaMA (Hierarchical/Causal Flow). $\mathcal{V}_{\text{tok}}$ comprises multi-scale motifs $\{v^{(s)}\}$, where the same molecule is re-represented across scales. $\psi_{ij}$ is determined by the causal mask $\mathbf{M}_{\text{causal}}$, enforcing a directed flow where coarse-scale tokens attend to fine-scale predecessors. This turns cross-scale aggregation into a learned, anisotropic connectivity along the sequence.
3.3.2 Information-Theoretic Context Analysis.
We contrast NTP with MNP. Let $x_{t}$ be a target token and let $Z_{t}$ denote its uncorrupted evidence set (e.g., the causal history for CamS coarse-scale tokens). MNP observes a stochastically corrupted context $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$ via random masking.
@@
For any non-trivial masking channel $\mathcal{M}$, we have
$I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t}).$ (12)
Sketch. The inequality follows from the Data Processing Inequality (DPI): masking is a stochastic post-processing of $Z_{t}$ that cannot increase mutual information (Cover, 1999). In graph MNP, this loss of information is exacerbated by evidence-pattern uncertainty: the most relevant local-neighborhood evidence for a masked token may be co-masked, making the conditional distribution unstable across masking patterns. Recent graph pretraining works explicitly mitigate this by masking structured units (e.g., triplets / local subgraphs) or designing structure-guided masking, and by injecting global knowledge tokens (Li et al., 2023; Rong et al., 2020; Liu et al., 2024). In contrast, CamS NTP for coarse-scale tokens conditions on an uncorrupted fine-to-coarse history, avoiding such stochastic evidence loss. This perspective is consistent with classic critiques of Masked Language Modelling (MLM): due to input corruption and factorized prediction over masked positions, it introduces a pre-train–fine-tune discrepancy and neglects dependencies among masked targets (Yang et al., 2019).
3.3.3 Gradient Flow Density.
We define the Direct Supervision Density (SD) as the expected fraction of tokens serving as regression targets per update. For a multi-scale sequence with total length $T_{\text{all}}=\sum_{s=1}^{S}T_{s}$ and mask ratio $\rho$:
$\text{SD}_{\text{NTP}}\approx 1\quad\gg\quad\text{SD}_{\text{MNP}}=\rho.$ (13)
This trade-off (more predicted tokens vs. more corruption) is central to masked modeling (Wettig et al., 2023). Empirically, KPGT finds the best masking ratio can be as high as $\rho=0.5$ when using triplet-style tokens and global knowledge nodes to stabilize context (Li et al., 2023). Meanwhile, dense token-level learning signals are known to substantially improve sample efficiency compared to sparse masked-only supervision (Clark et al., 2020). CamS further strengthens this effect by providing multi-view supervision: the same molecule contributes token-level targets repeatedly across scales.
3.3.4 Hierarchical Inductive Bias.
Graphformers rely on implicit depth and static structural biases to propagate information. In contrast, CamS injects an explicit structural hierarchy via the multi-scale sequence $X=[X_{1},\dots,X_{T_{\text{all}}}]$, and the causal mask ensures coarse-scale predictions have deterministic access to fine-scale details. We hypothesize this is particularly beneficial for activity-cliff regimes, where small local motif edits trigger global property shifts, as it biases the model toward bottom-up composition rules.
@@ 4.1 Experimental Setup
We decouple the pipeline into three stages, using the ChEMBL-34 Dataset for tokenizer training, the Enamine675M Dataset for NTP pre-training, and the MoleculeNet/MoleculeACE Benchmark for downstream fine-tuning and evaluation.
CamS-Tokenizer Training. CamS-tokenizer training is expensive because replaying $\mathcal{O}$ to materialize $\mathcal{V}$ repeatedly invokes RDKit’s subgraph-isomorphism matching (the dominant bottleneck) (Landrum et al., 2021; Ehrlich and Rarey, 2011) to extract motifs and attachment patterns, which is difficult to accelerate. We therefore train the tokenizer once on ChEMBL-34 (Gaulton et al., 2017), a moderately sized yet chemically diverse corpus, at the largest scale (67K). Concretely, we mine an ordered merge list $\mathcal{O}=(o_{1},\dots,o_{K})$ with K=685K{=}685 merges, apply it to obtain $\mathcal{V}$, and then perform SAVC to ensure single-atom coverage for common elements. Smaller scales are obtained without retraining by prefix slicing: each scale uses a prefix $\mathcal{O}_{\leq k}$ together with its induced vocabulary subset (i.e., motifs mined by applying $\mathcal{O}_{\leq k}$), where 1K uses k=0k{=}0, 7K/27K use k=62/210k{=}62/210, and 67K uses k=685k{=}685. Implementation details and full procedures are provided in Appendix A.1.
Pre-training Dataset and Configuration. We instantiate a 16-layer LLaMA-style decoder (hidden size 720, 8 attention heads) using HuggingFace Transformers. For self-supervised pre-training, we use Enamine675M (675M molecules) (Shivanyuk et al., 2007) and optimize NTP (Section 3.2.1). For each molecule, we construct five CamS training sequences: four single-scale BFS sequences tokenized with 1K/7K/27K/67K vocabularies, plus the concatenated multi-scale sequence in Eq. (1). We treat each sequence as one training instance and uniformly shuffle all instances, resulting in roughly $5\times$ as many sequences as molecules. All optimization details (schedule, batch size, steps) are summarized in Appendix C.1.
Downstream Benchmarks and Baselines. We evaluate on MoleculeNet (11 tasks; broad endpoints that probe general-purpose molecular property prediction) (Wu et al., 2018) and MoleculeACE (30 activity-cliff tasks; stress-testing fine-grained structural sensitivity under subtle edits that induce large activity shifts) (Van Tilborg et al., 2022). To enable an apples-to-apples comparison against a recent strong pre-training baseline, we adopt the data splits and evaluation protocol used in KPGT (Li et al., 2023). On MoleculeNet, following KPGT’s setup, we compare primarily against self-supervised pre-training baselines to isolate the effect of (1) representation/tokenization and (2) the pre-training objective. On MoleculeACE, we report the benchmark’s comprehensive ML/DL baselines, since activity cliffs emphasize discriminating small structural changes where strong feature-based methods remain competitive. Unless stated otherwise, CamS-LLaMA is kept at $\sim$100M parameters (comparable to KPGT); fingerprints are injected only during fine-tuning as a minimal prior, while pre-training is purely sequence-based.
4.2 Main Results
We summarize the main comparisons in Table 1 (MoleculeNet) and Table 2 (MoleculeACE).
General property prediction (MoleculeNet). Table 1 shows that CamS-LLaMA achieves overall state-of-the-art on MoleculeNet under a comparable model scale to KPGT ($\sim$100M): it slightly improves the average AUROC in classification (0.845 vs. 0.843) and the average RMSE in regression (1.172 vs. 1.175). We also win 6/11 individual tasks (5/8 classification and 1/3 regression), and improve over the previous best self-supervised learning baselines by a clear margin (AVG-AUROC: 0.845 vs. 0.825; AVG-RMSE: 1.172 vs. 1.272). Notably, this is achieved with weaker prior injection: unlike KPGT, which explicitly feeds rich handcrafted descriptors, we keep pre-training purely sequence-based and inject only a lightweight fingerprint prior during downstream fine-tuning. Full per-baseline results and extended tables are provided in Appendix D.

### v3

@@ 3.2 CamS-LLaMA
Autoregressive Pre-training via NTP. CamS-LLaMA adopts a standard LLaMA-style decoder (Touvron et al., 2023). In pre-training, CamS token lists are treated as ordinary sequences in the shared vocabulary $\Sigma$ and consumed under the usual causal mask. Concretely, to maximize structural learning signals via data augmentation, each molecule yields five sequence views (four single-scale + one multi-scale); we treat each view as an independent NTP training instance (details in Appendix A.3). Conditioned on the prefix context reshaped by fine-to-coarse concatenation (Eq. (1)), the model learns to predict the next token $x^{\star}_{t+1}$. The NTP loss is minimized on the pre-training corpus $\mathcal{D}_{\mathrm{pre}}$:
$\mathcal{L}_{\mathrm{NTP}}(\theta)=-\mathbb{E}_{X\sim\mathcal{D}_{\mathrm{pre}}}\sum_{t\in\mathcal{I}_{\mathrm{tok}}}\log p_{\theta}(x^{\star}_{t+1}\mid x_{\leq t}).$ (2)
Crucially, because coarse-scale tokens appear after fine-scale ones in $X$, prediction errors propagate gradients through the entire fine-scale prefix, effectively providing a fine-to-coarse causal credit assignment that embeds bottom-up compositional logic into the standard decoder.
Fine-tuning with Fingerprint Injection. Many methods inject rich handcrafted descriptors early during pre-training (Li et al., 2023). In contrast, we keep CamS as the primary representation and inject a lightweight fingerprint prior only during fine-tuning via a dual-path strategy (Early Injection + Late Fusion). Given a molecule’s fingerprint $\mathbf{f}\in\mathbb{R}^{D_{\mathrm{fp}}}$, we project it to $\mathbf{f}^{\prime}=W_{\mathrm{fp}}\mathbf{f}$ using a learnable projection $W_{\mathrm{fp}}\in\mathbb{R}^{H\times D_{\mathrm{fp}}}$. First, via Early Injection, $\mathbf{f}^{\prime}$ is prepended to the input embeddings as a global prompt:
$\mathbf{H}^{(0)}_{\mathrm{ft}}=[\,\mathbf{f}^{\prime};\mathbf{e}_{\text{BOS}};\dots;\mathbf{e}_{\text{EOS}}\,]\in\mathbb{R}^{(T+2)\times H}.$ (3)
After $L$ layers, we extract the representation of the [EOS] token, $\mathbf{h}_{\text{EOS}}^{(L)}$, which aggregates the AR structural context. Second, to allow the backbone to focus on deep implicit reasoning (by offloading explicit feature extraction to the direct path), we perform Late Fusion by concatenating $\mathbf{h}_{\text{EOS}}^{(L)}$ with the original projected fingerprint $\mathbf{f}^{\prime}$ to form the final representation $\mathbf{u}$:
$\mathbf{u}=[\,\mathbf{h}_{\text{EOS}}^{(L)}\,;\,\mathbf{f}^{\prime}\,]\in\mathbb{R}^{2H}.$ (4)
This shortcut not only preserves chemical fidelity but also serves as a stabilizer for fine-grained decision making (as analyzed in Section 4.4), ensuring predictions remain grounded even when deep features fluctuate. This fused vector is fed into task-specific heads. For classification ($\hat{\mathbf{y}}$) and regression ($\hat{y}$), the objectives are:
$\mathcal{L}_{\mathrm{cls}}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\big(\mathrm{softmax}(W_{2}\,\sigma(W_{1}\mathbf{u}))\big)_{y^{(i)}},$ (5)
$\mathcal{L}_{\mathrm{reg}}$ $=\frac{1}{B}\sum_{i=1}^{B}\big(W_{2}\,\sigma(W_{1}\mathbf{u})-y^{(i)}\big)^{2}.$
This strategy balances the pre-trained structural knowledge with the explicit fingerprint prior, optimizing $\mathbb{E}_{\mathcal{D}_{\mathrm{task}}}[\mathcal{L}_{\mathrm{task}}]$.
3.3 CamS LLaMA vs. Graph Transformer
@@
We analyze CamS-LLaMA and Graph Transformer-style models (e.g., KPGT) within a unified framework of Token-Level Graph-Structured Deep Feature Construction. We show that CamS yields (1) a denser multi-view direct supervision signal and (2) a hierarchical inductive bias that are mathematically distinct from standard masked-node prediction (MNP). Detailed derivations are in Appendix B.
@@
Unified Framework. Let $G=(V,E)$ be a molecular graph. Both paradigms construct a token graph $\mathcal{G}_{\mathrm{tok}}=(\mathcal{V}_{\mathrm{tok}},\mathcal{E}_{\mathrm{tok}})$ and perform $L$ layers of propagation:
@@
where $\mathbf{H}^{(l)}$ are token embeddings and $\psi_{ij}$ encodes structural bias. For Graph Transformers (Flat/Static Bias), $\mathcal{V}_{\mathrm{tok}}$ corresponds to single-scale atoms or triplets. $\psi_{ij}$ encodes static biases (e.g., shortest-path distance) independent of token content, yielding an isotropic receptive field (Ying et al., 2021). For CamS-LLaMA (Hierarchical/Causal Flow), $\mathcal{V}_{\mathrm{tok}}$ comprises multi-scale connection-aware motifs $\{v^{(s)}\}$. $\psi_{ij}$ is determined by the causal mask $\mathbf{M}_{\mathrm{causal}}$, enforcing a directed flow where coarse-scale tokens attend to fine-scale predecessors (via Inter-Scale Order). This turns cross-scale aggregation into a learned, anisotropic connectivity.
Information-Theoretic Context Analysis. We contrast NTP with MNP. Let $x_{t}$ be a target token and $Z_{t}$ its uncorrupted evidence set. MNP observes a stochastically corrupted context $\tilde{Z}_{t}=\mathcal{M}(Z_{t})$.
@@
For any non-trivial masking channel $\mathcal{M}$, we have $I(x_{t};Z_{t})\geq I(x_{t};\tilde{Z}_{t})$.
Sketch. The inequality follows from the Data Processing Inequality (DPI): masking is a stochastic post-processing of $Z_{t}$ that cannot increase mutual information (Cover, 1999). In graph MNP, this loss of information is exacerbated by evidence-pattern uncertainty: the most relevant local-neighborhood evidence for a masked token may be co-masked, making the conditional distribution unstable across masking patterns. Recent graph pretraining works explicitly mitigate this by masking structured units (e.g., triplets or local subgraphs), designing structure-guided masking, or injecting global knowledge tokens (Li et al., 2023; Rong et al., 2020; Liu et al., 2024). In contrast, CamS NTP for coarse-scale tokens conditions on an uncorrupted fine-to-coarse history, avoiding such stochastic evidence loss. This perspective is consistent with classic critiques of Masked Language Modelling (MLM): due to input corruption and factorized prediction over masked positions, it introduces a pre-train–fine-tune discrepancy and neglects dependencies among masked targets (Yang et al., 2019).
Gradient Flow Density and Hierarchical Inductive Bias. We define the Direct Supervision Density (SD) as the expected fraction of tokens serving as targets per update. With mask ratio $\rho$, we have $\text{SD}_{\text{NTP}}\approx 1\gg\text{SD}_{\text{MNP}}=\rho$. Masked models face a dilemma: increasing $\rho$ improves density but exacerbates context corruption (Wettig et al., 2023); for instance, KPGT relies on global nodes to stabilize optimization at $\rho=0.5$ (Li et al., 2023). In contrast, CamS achieves maximal density on uncorrupted contexts and the same structure contributes gradients repeatedly across scales, yielding substantially higher sample efficiency than sparse masked supervision (Clark et al., 2020). From another perspective, Graph Transformers rely on implicit depth to propagate information. In contrast, CamS injects an explicit structural hierarchy via the multi-scale sequence $\mathbf{X}$, where the causal mask ensures coarse-scale predictions have deterministic access to fine-scale details. We hypothesize this bias is particularly beneficial for stabilizing predictions in activity-cliff regimes (as supported by ablation study in Section 4.4).
@@ 4.1 Experimental Setup
CamS-Tokenizer Configuration. We train the CamS-Tokenizer on ChEMBL-34 (Gaulton et al., 2017) to balance diversity with feasibility, as vocabulary materialization involves expensive subgraph-isomorphism matching (Landrum et al., 2021; Ehrlich and Rarey, 2011) that prohibits direct training on billion-scale corpora. We learn the merge list $\mathcal{O}$ once with K=685K{=}685 operations to materialize the full CamS Vocabulary ($67\mathrm{K}$). To instantiate the multi-scale context (M=4M{=}4), we slice $\mathcal{O}$ at prefix indices $k\in\{0,62,210,685\}$, inducing Motif Scales $s\in\{1\mathrm{K},7\mathrm{K},27\mathrm{K},67\mathrm{K}\}$. These scales are selected to span the granularity spectrum: the $1\mathrm{K}$ scale serves as the atomic baseline (fine limit), the $67\mathrm{K}$ scale represents the experiment-maximal coarse limit, while $7\mathrm{K}$ and $27\mathrm{K}$ provide necessary intermediate resolutions to bridge the gap. Full details are in Appendix A.1.
Pre-training Implementation. We instantiate a 16-layer LLaMA decoder (hidden size 720, 8 heads; $\sim$100M parameters, comparable to KPGT). Pre-training uses Enamine675M (675M molecules) (Shivanyuk et al., 2007). Following Section 3.2, we generate five sequence views per molecule, treating them as independent instances and uniformly shuffling the resulting dataset. We optimize the NTP objective (Eq. 2) using hyperparameters summarized in Appendix C.1.
Downstream Benchmarks and Baselines. We evaluate on MoleculeNet (11 general property tasks) (Wu et al., 2018) and MoleculeACE (30 activity-cliff tasks) (Van Tilborg et al., 2022), strictly adhering to the data splits and protocols established by the SOTA baseline KPGT (Li et al., 2023) (details in Appendix C.2). We benchmark against KPGT and the baselines adopted in KPGT’s evaluation. Crucially, distinct from KPGT’s heavy reliance on rich descriptor at all stages, we inject lightweight fingerprints only during fine-tuning, maintaining a purely structural pre-training. On MoleculeNet, we include comprehensive graph-based self-supervised baselines, such as MolFormer (MolF) (Wu et al., 2023), ContextPred (Ctxt) (Hu et al., 2020), GROVER (GROV) (Rong et al., 2020), JOAO (You et al., 2020), GEM (Fang et al., 2022), GraphMAE (GMAE) (Hou et al., 2022), and MoleBERT (MBRT) (Xia et al., 2023). Additionally, we benchmark against the SMILES-NTP FM NatureLM (Xia et al., 2025) (comparison in Appendix D). On MoleculeACE, we additionally include strong classical machine learning baselines (SVM, RF, GBM (Van Tilborg et al., 2022) with ECFP (E) (Rogers and Hahn, 2010) or MACCS (M) (Durant et al., 2002) descriptors), which are known to outperform deep learning methods on the activity-cliff stress test.
4.2 Results on Downstream Tasks
General Property Prediction (MoleculeNet). Table 1 demonstrates that CamS-LLaMA establishes a new SOTA on MoleculeNet. While maintaining a comparable model scale to KPGT ($\sim$100M) with fewer priors, CamS-LLaMA surpasses it in both classification (AVG-AUROC: 0.845 vs. 0.843) and regression (AVG-RMSE: 1.172 vs. 1.175). Crucially, these results underscore superior scalability and efficiency: unlike KPGT’s descriptor-imposed bottleneck on data scaling or NatureLM’s reliance on massive parameters ($\sim$56B, see Appendix D), CamS leverages a purely sequence-driven pre-training to efficiently scale to 675M pure molecular data (discussion in Appendix E) and achieve SOTA with a compact backbone. By winning 6 out of 11 individual tasks, CamS-LLaMA proves that properly serialized causal modeling captures molecular properties as effectively as descriptor-enhanced graph methods, without their reliance on external domain priors. Full per-baseline results are provided in Appendix F.